From: STOSC::SMM "On-Duty SMM x5074" 16-JAN-1993 01:51:34.20 To: @[.DIS]HST_STAT CC: Subj: HST Status Report, 01/14 - 01/15. HUBBLE SPACE TELESCOPE DAILY REPORT #804 PERIOD COVERED: 0800 EST 01/14/93 - 0800 EST 01/15/93 Daily Status Report as of 015/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC 2438 (Study of the Most Distant Galaxies) The Wide Field Camera was used to make six images of the high red shifted galaxy 4C41.17 using a red shifted Lyman Alpha and a continuum filter. All six images were received and displayed. The first image showed one packet of data missing. The Data Capture Facility (DCF) reported that the frame sync patterns were off by as much a six bits. The packet was reported to be unrecoverable, and the problem could be due to transmission errors anywhere downstream of the TDRSS satellite. All of the other images were received and displayed. They were reported to look okay. 1.2 Completed GHRS/2 3755 (Absorption Cross Sections of Nearby Galaxies) The High Resolution Spectrograph was used to observe quasi-stellar object 0637-752. This object was used as a probe to study the absorption lines produced by the halo of a foreground galaxy. The return to brightness spiral search was executed in course track with no problems, and the target was found in the sixth map. A status buffer error message was received during an LOS which indicated that there had been a failure to position the carousel. An HSTAR was written. The observation was restarted after the reacquisition, and went well. In a subsequent observation one packet of data was scrambled. Preliminary analysis indicated that this is the same problem which was reported in section 1.1 . The other observations were reported to be okay. 1.3 Completed PC 4714 (Internal Delta Flats) The Planetary Camera was used to make a single internal flat observation. This image will be used for the restoration of Earth flats, which were altered due to changes in the quantum efficiency which occurred after WFPC decontamination operations. The observations was reported to be okay. 1.4 Completed FOS/BL 4110 (The Physics of Massive O-type Stars) The Faint Object Spectrograph was used to make observations of OB Class star HD141556 which is located in the Large Magellanic Cloud. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 14 Successful : 14 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 5 Successful : 5 2.3 Significant HSTARs: None 2.4 Operations Note: The third Servicing Mission Internal Simulation (aka. EVA Sim. #1) was completed yesterday (Wednesday, Jan.14) from 014/1300 through 014/2300Z. This simulation covered the time line from the Solar Array Primary Deployment Motor (PDM) event through the conclusion of EVA #1. This simulation provided the HST Servicing Mission team with functional training to ensure operational readiness. The test went well, and no major problems were reported. Commands were sent to the PSEA configuration memory to revert back to the original MSS configuration. The MSS reasonableness test has been reactivated so the PSEA will automatically select MSS 2 if the MSS 1 signals are out of limits. Since both MSSs have exhibited anomalous signatures at times, there is no longer a strong reason to pre-select the MSS to be used. The new configuration gives the PSEA maximum flexibility in selection an MSS. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Three Sets of WFC 1297 (GTO High Latitude Parallel Program) 3.2 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.3 Complete GHRS/2 3372 (Point Spread Function Determination) 3.4 Complete PC 4140 (Rapid Internal Monitor - Part 2) 3.5 Complete FOS/BL 4057 (Spectrography of QSOs) 3.6 Complete FOS/BL 4110 (The Physics of Massive O-Type Stars) 3.7 Complete WFC-PC 4109 (Rapid Internal Monitor) 3.8 Complete Two Sets of PC 4714 (Rapid Internal Monitor) 3.9 Complete FOS/RD/BL 3647 (Star forming Histories of Elliptical Galaxies) 3.10 Complete WFC 4105 (HST Medium Deep Survey High Latitude) 3.11 Complete Two Sets of WFC 3917 (HST Medium Deep Survey - Part 1) 3.12 Complete FOC/96 3880 (Absorption Lines in Planetary Nebulae) 3.13 Complete WFC 1306 (GTO Low Latitude Parallel Program) 3.14 Complete FOS 4000 (Quasar Absorption Line Survey) 3.15 Complete FOC 3801 (Search for QSOs Suitable for Subsequent Observations) 3.16 Complete FOC 3815 (UV Observations of Hubble Sandage Variables) 3.17 Complete PC 4764 (Investigation of Dynamics of Binary Asteroids) 3.18 Complete Three Sets of FOS/RD 3744 (Spectrophotometry of Phobos and Demos) 3.19 Complete PC 3603 (Snapshots of Protoplanetary Nebulae) 3.20 Complete FOS/RD/BL 4694 (Stellar Content of Galaxies and Globular Clusters) 3.21 Complete PC 4106 (HST Medium Deep Survey: Low Latitude) 3.22 Complete PC 4712 (Measles Monitor) 3.23 Complete PC 3230 (PSF and Focus Monitoring 2) 3.24 Complete HSP/UV2 3798 (Search for Radial Pulsations in White Dwarfs) 3.25 Complete FOS/BL 3816 (White Dwarf Stars) 3.26 Initiate HSP/UV2 1095 (Variability of High Luminosity Stars) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-